OpenStudy (anonymous):

-1/4 / 5 help please! / = divided

OpenStudy (anonymous):

please write full equation

OpenStudy (unklerhaukus):

\[\frac{-1/4}{5}\text { or }\frac{-1}{4/5}\]

OpenStudy (lgbasallote):

\[\huge \frac{-\frac 14}{5} \implies \frac{-\frac 14}{\frac 51} \implies -\frac 14 \times \frac 15\]

OpenStudy (anonymous):